OpenStudy (anonymous):

evaluate [(30+6)-3²]/9*2

OpenStudy (anonymous):

6?

thomaster (thomaster):

you need to simplify it?

OpenStudy (anonymous):

NO

thomaster (thomaster):

\(\huge\frac{30+6-3^2}{9*2}\)

OpenStudy (anonymous):

3/2

thomaster (thomaster):

@aras_birdal Don't just give an answer. Guide the asker to the answer. Read the \(\bf\color{red}{\href{/code-of-conduct}{Code~of~Conduct}}\)

OpenStudy (anonymous):

36 MINUS 9 IS 27, THEN SIMPLIFY 27 AND 9 THERE IS LEFT 3 ON TOP AND THERE WAS 2 AT BELOW SO THE ANSWER IS 3/2

thomaster (thomaster):

That's correct yea, no need to write in caps though \(\Huge\ddot\smile\) @demick Do you understand how to get to the answer?

OpenStudy (anonymous):

yes and please just give me the answer

thomaster (thomaster):

No we don't give answers here. Only guidance. \(\huge\frac{30+6-3^2}{9*2}\) \(\huge\frac{36-3^2}{9*2}~\to~\huge\frac{36-9}{9*2}~\to~\huge\frac{27}{9*2}\) Do you know what to do next?

thomaster (thomaster):

@demick